OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#4453 — Optic HS th1-1-
Incident Report for Network & Infrastructure
Resolved
Regarding the links' migration (Roubaix's side) task #4440, one of the optics has been disconnected at Paris's side:

Optical Optical
Temperature Voltage Current Tx Power Rx Power
Port (Celsius) (Volts) (mA) (dBm) (dBm)
------- ----------- ------- -------- -------- --------
Te5/2 31.7 0.00 1.1 -- -13.6 -- -2.2

More power output. A technician intervened to replace by a spare.


Update(s):

Date: 2010-08-11 00:30:15 UTC
The optic has been replaced around 12:30.
Posted Aug 11, 2010 - 00:27 UTC